OpenStudy (anonymous):

Precalc, please help? Find the distance between the points with polar coordinates (3, 120°) and (0.5, 49°). I got 2.035, but I am not sure if that is right.

OpenStudy (anonymous):

x=cos theta y=rsin theta. \[x1=r \cos \theta =3\cos 120=3\cos \left( 180-60 \right)=-3\cos 60=\frac{ -3 }{ 2 }\] \[y1=rsin \theta=3\sin 120=3\sin \left( 180-60 \right)=3\sin 60=3*\frac{ \sqrt{3} }{ 2 },\] similarly find x2,y2 then \[d=\sqrt{\left( x2-x1 \right)^2+\left( y2-y1 \right)^2}\]